16 Jordan Company manufactures molded candles that are finished by hand. The company developed the following standards for aneed help with all parts pls
Answer #1

a&b.

Cost Standard cost per unit
Direct material $1.08 (1.80*$0.60)
Direct labor 8.01 (0.90*$8.90)
Overhead 4.7 ($159,800/34,000)
Total per candle $13.79

c&d.

Cost Actual cost per unit
Direct material $1.04 (48,300*$0.56 / 26,000)
Direct labor 8.26 (26,200*$8.20 / 26,000)
Overhead 4.9 ($127,400/26,000)
Total per candle $14.2

e.

Direct material price variance = Actual quantity * (Standard price-Actual price)

Direct material price variance = 48,300 * $0.04 ($0.60-0.56) = $1,932 Favorable

Direct material usage variance = Standard price * (Standard quantity-Actual quantity)

Standard quantity = 26,000*1.80 = 46,800

Direct material usage variance = $0.60 * (46,800-48,300) = $900 Unfavorable

Direct labor price variance = Actual hours * (Standard price-Actual price)

Direct labor price variance = 26,200 * ($8.90-8.20) = $18,340 Favorable

Direct labor usage variance = Standard price * (Standard hours-Actual hours)

Standard hours = 26,000*0.90 = 23,400 hours

Direct labor usage variance = $8.90 * (23,400-26,200) = $24,920 Unfavorable

f.

Fixed overhead spending variance = Actual fixed overhead - Budgeted fixed overhead

Fixed overhead spending variance = $127,400 - 159,800 = $32,400 Favorable

Fixed overhead volume variance = Budgeted fixed overhead - Applied fixed overhead

Fixed overhead volume variance = $159,800 - 122,200 (26,000*$4.7) = $37,600 Unfavorable